# Breaking through the learning plateaus of in-context learning in Transformer

### v2

@@ 6 Conclusion
This paper establishes a connection between the weights components and learning plateaus. Building on this connection, the paper proposes three strategies to overcome these plateaus. These strategies have proven to be effective in both synthetic and natural language tasks.
Author Constribution
J.Fu lead the Research with the support from Y.Wang, Y.Lu and N.Zheng. The conceptualization of the study was orchestrated by J.Fu, who was aided by Y.Wang. Y.Tao pointed out that Shape3D can be used for exploring in-context learning and designed the synthetic task based on Shape3D (Fig. 2). The experimental work was executed by J.Fu, who also benefited from Y.Tao’s contribution in preparing the Shape3D dataset. J.Fu took the lead in drafting the manuscript, with W.Yu providing support. W.Yu was responsible for the creation of Fig. 2 and Fig. 7, while J.Fu produced the remaining figures. The proof was undertaken by J.Fu, with Y.Tao offering assistance in verifying the accuracy of the proofs. All the authors participated in the discussion of this paper.
@@ Appendix A Detail of experiments on sythetic tasks
3dshapes [^1] https://github.com/deepmind/3d-shapes [22] is a dataset of 3D shapes procedurally generated from 6 ground truth independent latent factors. These factors are floor colour, wall colour, object colour, scale, shape and orientation. All possible combinations of these latents are present exactly once, generating N = 480000 total images. Latent factor values including: 1) floor hue: 10 values linearly spaced in [0, 1], 2) wall hue: 10 values linearly spaced in [0, 1], 3) object hue: 10 values linearly spaced in [0, 1], 4) scale: 8 values linearly spaced in [0, 1], 5) shape: 4 values in [0, 1, 2, 3], and 6) orientation: 15 values linearly spaced in [-30, 30].
@@
We employ the Adam optimizer [23] and mini-batch training to optimize the loss function $\mathcal{L}(\theta,S)$. Here, we use cross-entropy as the loss function. We utilize a batch size of 128 and set the learning rate to 0.0001. For training purposes, we use $10^{5}$ sequences and, for evaluation, $4\times 10^{4}$ sequences. There is no overlap between images in the training sequences and those in the evaluation sequences.
@@ A.3 More detail about probe framework
We employ metrics for numerical evaluation of components and in-context learning performance. Since the components are hidded in the representation, we use the probe method [3]. The probe classifier has a single linear layer, with softmax and cross-entropy calculating the loss. Because we don’t know the specific form of $g_{weights}$ and $g_{context}$, we choose the representation within the Transformer that given the high score for weights and context components for approximating. We use the linear probe because using more complex probe model doesn’t have significant improve in accuracy and the change of the accuracy during the training process is more important than the absolutely value. The probe model is trained utill totally converge. The probe is trained for 2 epoch for context component and 1 epoch for weights component. The in-weight probe predicts values of six factors of all images, while the in-context probe identifies the hidden factor for each sequence. The details are as follows.
@@ B.3 Dataset Detail
SST-ICL dataset  The dataset is contructed based on SST [41] datasets. We remove the long review in the datasets and transform the original labels into “Negative", “Positive" and “Neutral". Then, we organize the reviews follow the same way as that in Subsection A.4. We produce $10^{4}$ sequence for training and $4\times 10^{3}$ for testing. Each sequence contains 5 reviews. We illustrate the example of the dataset in Fig. 6 AB.
@@ C.1 In-weights and in-context learning
Previous studies [10, 11] examined the relationship between in-weights learning and in-context learning. The division of in-weights learning and the in-context learning process is conceptually similar to our distinction between the influence of the weights and context component. In in-context learning, the Transformer relies on a combination of both its internal weights and the context provided to address a given task. In contrast, for in-weights learning, the model depends exclusively on its weights. Nevertheless, there are notable differences between these two notions: 1) The internal weights and context components are concerned with capturing information within the Transformer, whereas the focus of in-weights and in-context learning is on how the Transformer tackles a task. In-weights learning is a rename for regular supervised learning. 2) Both weights and context components coexist and play roles within the paradigm of in-context learning. While in-weights learning is distinguished from in-context learning.
@@
1. Min et al. [32] discovered that (1) performance can be improved by increasing the number of in-context examples. (2) Changing the labels of in-context examples does not influence the predicted label. The first discovery indicates that the prediction relies on the context components. The second discovery suggests that the Transformer uses the weights component for label prediction, given that there is no observed change when the labels of in-context examples are altered.
2. Brown et al. [8] found that larger models are increasingly effective at utilizing in-context information. This suggests that in real-world scenarios, the efficiency of the context component improves with the enlargement of the model’s size. Brown et al. [8] also found that enhancing the model size can boost its zero-shot capabilities. These findings suggest that scaling the model can enhance both the weights and context components, and the model employs these two components to address the problem.
3.[48] carried out research on a two-class classification issue. They conducted experiments in which they altered a certain percentage of labels in the context examples to ascertain if the model’s prediction would also change. If a change was observed, it would imply that the prediction relies on the context components. If no change was noticed, the prediction would be considered to depend on the weights component. Their results were intermediate, suggesting that both weights and context components contribute. Additionally, they found that enhancing the model size increases the impact of in-context examples.
@@ C.3 Related works for understanding Transformer
Analysis of Transformer  The analysis of Transformers can be broken down into two main components: examining the expressibility of Transformers and comprehending the mechanisms of learned Transformers. To analyze the expressibility of Transformers, a common approach is to determine if they can solve specific problems by constructing appropriate weights. Giannou et al. [19] demonstrates that Transformers can function as Turing machines, while Liu et al. [28] shows that they can learn shortcuts to solve automata problems. In addition to expressibility, researchers have also investigated the mechanisms behind learned Transformers. Bietti et al. [7] examines Transformers from a memory standpoint, and Tian et al. [42] focuses on single-layer Transformers. While the analysis of Transformers is crucial to our work, our ultimate goal differs; we aim to bridge the gap between representation learning and in-context learning.
Exploration of representation within Transformer.  Owing to the widespread use of Transformers, numerous studies [26, 44] seek to investigate their internal representations as a means of comprehending their functionality. The most prevalent approach involves utilizing probe models and tasks to discern the information stored within these representations [44, 37]. Taking a different perspective, Voita et al. [45] explores the flow of information across Transformer layers and how this process is influenced by the selection of learning objectives. Our work shares similarities with these studies in that we employ the probe method to examine representations. However, our focus differs in that we do not concentrate on the semantic meaning within the representation. Instead, we investigate how the weights and in-context information impact representation.
@@ Appendix E Contruction analysis of why the weights component is related to the learning plateaus
In our analysis, we explore a modified, more relaxed variant of the Transformer model. The rationale behind this relaxation is underpinned by evidence suggesting that 1) employing the ReLU activation function in the feed-forward layers can achieve results comparable to the original model[34], and 2) the softmax operation may not be essential for the functioning of the Transformer[49, 9, 36]. The relaxed Transformer is defined as follows:
@@
(Transformer) One layer of Transformer contains one attention layer and one MLP layer. The calculation of Attention Layer is
@@
[…]
The constucted Transformer achieve significant performance.  To illustrate that the construction is meanningful, we compare the performance of the constructed model with the previous trained model. We choose the performance of the model trained on $D_{\text{fix}\land\text{rnd}}$ settings as it is the best performed model on $D_{\text{rnd}}$ setting. We find that the constructed model achieved a comparable performance as that trained model. This results indicate that our construction is meaningful.
@@ E.1 Proof Sketch
1. Estimate the hidden factor.  According to the perfect weights component assumption, for any $j$, we can project the $j$-th token feature into the space ${\mathbf{f}}^{(e)}_{j}$. Assume the $j$-th token is not the prompt token, then we have its label information $y_{j}$. Then, for $i$-th token, where $i<j$, obviously $i$ is also not the prompt token. As the results, we also have the information about $y_{i}$. If a factor $e$ is the hidden factor, then we would expect ${\mathbf{y}}_{i}\cdot{\mathbf{y}}_{j}={\mathbf{m}}(v^{(e_{h})}_{i})\cdot{\mathbf{m}}(v^{(e_{h})}_{j})={\mathbf{v}}^{(e_{h})}_{i}\cdot{\mathbf{v}}^{(e_{h})}_{j}={\mathbf{f}}_{i}^{(e_{h})}\cdot{\mathbf{f}}_{j}^{(e_{h})}$, where ${\mathbf{v}},{\mathbf{y}},{\mathbf{m}}$ is the corresponding one-hot version of $v,y,m$. Therefore, if we can find $e$ such that ${\mathbf{f}}_{i}^{(e)}\cdot{\mathbf{f}}_{j}^{(e)}$ can have a same value as ${\mathbf{y}}_{i}\cdot{\mathbf{y}}_{j}$, for all $i$, then $e$ can be predicted as hidden factor. Based on this intuition, in the construction, we focus on find a way to compare the value between ${\mathbf{f}}_{i}^{(e)}\cdot{\mathbf{f}}_{j}^{(e)}$ and ${\mathbf{y}}_{i}\cdot{\mathbf{y}}_{j}$. All this operations are done in the first layer.
2. Block unrelated information.  After find the hidden factor, the next step is to block the information that unrelated to the hidden factor. Blocking the unrelated information can remove the influence of it and simplify the following steps. We add the large negative value to the positions of the representation that is unrelated to the hidden factor. Then, through the Relu operation, all these negative value will be removed. These operations are placed in Layer 2.
3. Predict $y$.  The final step to predict the $y_{p}$ for the prompt sample ${\mathbf{x}}_{p}$. The challenge here is that we don’t know the mapping function $m$ that bridge the factor value and the corresponding label. Consider the relation that for any $i,j$ satisfying $i<j$, we have ${\mathbf{v}}_{i}^{(e_{h})}\cdot{\mathbf{v}}_{j}^{(e_{h})}={\mathbf{f}}_{i}^{(e_{h})}\cdot{\mathbf{f}}_{j}^{(e_{h})}={\mathbf{y}}_{i}\cdot{\mathbf{y}}_{j}$. Therefore, if we can find a $i$, satisfying that ${\mathbf{f}}_{i}^{(e_{h})}\cdot{\mathbf{f}}_{j}^{(e_{h})}=1$, then we have $y_{i}=y_{j}$. Then, we can copy $y_{i}$ to the representation of $j$-th token and use the final linear layer to output $y_{i}$ as the prediction for $y_{j}$.
@@
Given the perfect weights component assumption, there are two source that cause the error prediction. We give the seperate analyzing below.
1. Fail to find the correct hidden factor. Finding correct hidden factor is essential to make correct prediction. The correct hidden factor cannot be infered if there are more than two factors such that the values of ${\mathbf{f}}_{i}^{(e)}\cdot{\mathbf{f}}_{j}^{(e)}$ and ${\mathbf{y}}_{i}\cdot{\mathbf{y}}_{j}$ can be matched according to the construction of first layer.
2. Make error prediction give correct factor. Given hidden factor, it is also possible that we cannot make correct prediction if we fail to associate the correct label to the factor value. This will happen if all the context examples don’t contain a same factor value of hidden factor as the prompt image.
@@ E.3 Construction of Transformer
Without loss of generality, we assume the representation of Transformer in layer $k$ is in a form that ${\mathbf{h}}_{2i-1}^{(k)}=({\mathbf{f}}_{i},\mathbf{0},\mathbf{0},\mathbf{0},\mathbf{0},\mathbf{0},\mathbf{0},{\mathbf{p}}_{i})^{\mathrm{T}}$ and ${\mathbf{h}}_{2i}^{(k)}=(\mathbf{0},{\mathbf{y}}_{i},\mathbf{0},\mathbf{0},\mathbf{0},\mathbf{0},\mathbf{0},{\mathbf{p}}_{i})^{\mathrm{T}}$ (Remind that one sample will takes two token, one for $x$ and one for $y$). Because the representation usually lays in low dimension space, a simple linear layer can transfer the representation in our defined sparse form. What’s more, it is nature to assume that the position information is stored in the representation, since it is given in the input and it is essential for attention.
@@
This layer obtain the logit of new sample by comparing the similarity between the unblocked feature of this sample and the in-context sample.
@@
Note that value ${\mathbf{f}}_{i,x}^{\prime}\cdot{\mathbf{f}}_{a,y}^{\prime}$ is equal to the number of unblocked factors (both unbloked) that have same value between $a$-th sample and $i$-th sample Obviously, there is a $W_{O}^{(l+3,1)}$ such that ${\mathbf{h}}_{2i-1}=({\mathbf{f}}_{i,x}^{\prime},\mathbf{0},{\mathbf{u}}_{2i-1},\mathbf{0},\mathbf{0},\mathbf{0},{\mathbf{m}}_{2i-1}^{\prime},\text{Logit},{\mathbf{p}}_{i})$.

### v3

@@ 7 Conclusion
This paper establishes a connection between the weights component and learning plateaus. Building on this connection, the paper proposes three strategies to overcome these plateaus. These strategies have proven to be effective in both synthetic and natural language tasks.
8 Limitation
In this paper, we mainly focus on understanding of the learning plateaus of in-context learning with Transformer, where the Transformer requires the information from the context examples and the query example to make a prediction. Therefore, our method cannot explain the learning plateaus phenomenon outside this score. There are some works (Nanda et al., 2023; Power et al., 2022) that discuss the learning plateaus phenomenon in supervised learning. Our work fails to explain these phenomena.
Acknowledgements
Jingwen Fu, Tao Yang, and Nanning Zheng are supported by the National Natural Science Foundation of China (Grant No. 62088102).
Impact Statement
The goal of this paper is to understand the mechanism of learning plateaus in Transformers and find a method to avoid learning plateaus. Our work doesn’t have a direct influence on society. However, future works based on our work may influence society but it is unpredictable currently.
@@ Appendix A Detail of experiments on sythetic tasks
3dshapes [^1] https://github.com/deepmind/3d-shapes (Kim & Mnih, 2018) is a dataset of 3D shapes procedurally generated from 6 ground truth independent latent factors. These factors are floor colour, wall colour, object colour, scale, shape and orientation. All possible combinations of these latents are present exactly once, generating N = 480000 total images. Latent factor values including: 1) floor hue: 10 values linearly spaced in [0, 1], 2) wall hue: 10 values linearly spaced in [0, 1], 3) object hue: 10 values linearly spaced in [0, 1], 4) scale: 8 values linearly spaced in [0, 1], 5) shape: 4 values in [0, 1, 2, 3], and 6) orientation: 15 values linearly spaced in [-30, 30].
@@
We employ the Adam optimizer (Kingma & Ba, 2014) and mini-batch training to optimize the loss function $\mathcal{L}(\theta,S)$. Here, we use cross-entropy as the loss function. We utilize a batch size of 128 and set the learning rate to 0.0001. For training purposes, we use $10^{5}$ sequences and, for evaluation, $4\times 10^{4}$ sequences. There is no overlap between images in the training sequences and those in the evaluation sequences.
@@ A.3 More detail about probe framework
We employ metrics for numerical evaluation of components and in-context learning performance. Since the components are hidded in the representation, we use the probe method (Alain & Bengio, 2016). The probe classifier has a single linear layer, with softmax and cross-entropy calculating the loss. Because we don’t know the specific form of $g_{weights}$ and $g_{context}$, we choose the representation within the Transformer that given the high score for weights and context components for approximating. We use the linear probe because using more complex probe model doesn’t have significant improve in accuracy and the change of the accuracy during the training process is more important than the absolutely value. The probe model is trained utill totally converge. The probe is trained for 2 epoch for context component and 1 epoch for weights component. The in-weight probe predicts values of six factors of all images, while the in-context probe identifies the hidden factor for each sequence. The details are as follows.
@@ B.3 Dataset Detail
SST-ICL dataset  The dataset is contructed based on SST (Socher et al., 2013) datasets. We remove the long review in the datasets and transform the original labels into “Negative”, “Positive” and “Neutral”. Then, we organize the reviews follow the same way as that in Subsection A.4. We produce $10^{4}$ sequence for training and $4\times 10^{3}$ for testing. Each sequence contains 5 reviews. We illustrate the example of the dataset in Fig. 7 AB.
@@ C.1 In-weights and in-context learning
Previous studies (Chan et al., 2022a, b) examined the relationship between in-weights learning and in-context learning. The division of in-weights learning and the in-context learning process is conceptually similar to our distinction between the influence of the weights and context component. In in-context learning, the Transformer relies on a combination of both its internal weights and the context provided to address a given task. In contrast, for in-weights learning, the model depends exclusively on its weights. Nevertheless, there are notable differences between these two notions: 1) The internal weights and context components are concerned with capturing information within the Transformer, whereas the focus of in-weights and in-context learning is on how the Transformer tackles a task. In-weights learning is a rename for regular supervised learning. 2) Both weights and context components coexist and play roles within the paradigm of in-context learning. While in-weights learning is distinguished from in-context learning.
@@
1. Min et al. (2022) discovered that (1) performance can be improved by increasing the number of in-context examples. (2) Changing the labels of in-context examples does not influence the predicted label. The first discovery indicates that the prediction relies on the context components. The second discovery suggests that the Transformer uses the weights component for label prediction, given that there is no observed change when the labels of in-context examples are altered.
2. Brown et al. (2020) found that larger models are increasingly effective at utilizing in-context information. This suggests that in real-world scenarios, the efficiency of the context component improves with the enlargement of the model’s size. Brown et al. (2020) also found that enhancing the model size can boost its zero-shot capabilities. These findings suggest that scaling the model can enhance both the weights and context components, and the model employs these two components to address the problem.
3.(Wei et al., 2023) carried out research on a two-class classification issue. They conducted experiments in which they altered a certain percentage of labels in the context examples to ascertain if the model’s prediction would also change. If a change was observed, it would imply that the prediction relies on the context components. If no change was noticed, the prediction would be considered to depend on the weights component. Their results were intermediate, suggesting that both weights and context components contribute. Additionally, they found that enhancing the model size increases the impact of in-context examples.
@@ C.3 Related works for understanding Transformer
Analysis of Transformer  The analysis of Transformers can be broken down into two main components: examining the expressibility of Transformers and comprehending the mechanisms of learned Transformers. To analyze the expressibility of Transformers, a common approach is to determine if they can solve specific problems by constructing appropriate weights. Giannou et al. (2023) demonstrates that Transformers can function as Turing machines, while Liu et al. (2022) shows that they can learn shortcuts to solve automata problems. In addition to expressibility, researchers have also investigated the mechanisms behind learned Transformers. Bietti et al. (2023) examines Transformers from a memory standpoint, and Tian et al. (2023) focuses on single-layer Transformers. While the analysis of Transformers is crucial to our work, our ultimate goal differs; we aim to bridge the gap between representation learning and in-context learning.
Exploration of representation within Transformer.  Owing to the widespread use of Transformers, numerous studies (Li et al., 2022; Voita & Titov, 2020) seek to investigate their internal representations as a means of comprehending their functionality. The most prevalent approach involves utilizing probe models and tasks to discern the information stored within these representations (Voita & Titov, 2020; Schouten et al., 2022). Taking a different perspective, Voita et al. (2019) explores the flow of information across Transformer layers and how this process is influenced by the selection of learning objectives. Our work shares similarities with these studies in that we employ the probe method to examine representations. However, our focus differs in that we do not concentrate on the semantic meaning within the representation. Instead, we investigate how the weights and in-context information impact representation.
@@ Appendix E Contruction analysis of why the weights component is related to the learning plateaus
In our analysis, we explore a modified, more relaxed variant of the Transformer model. The rationale behind this relaxation is underpinned by evidence suggesting that 1) employing the ReLU activation function in the feed-forward layers can achieve results comparable to the original model (Press et al., 2019), and 2) the softmax operation may not be essential for the functioning of the Transformer (Wiegreffe & Pinter, 2019; Brunner et al., 2019; Richter & Wattenhofer, 2020). The relaxed Transformer is defined as follows:
@@
(Transformer) One layer of the Transformer contains one attention layer and one MLP layer. The calculation of the Attention Layer is
@@
We consider the data with $n_{e}$ factors and each factor has $n_{v}$ values in $D_{\text{rnd}}$ setting. For causal Transformer with the number of heads larger or equal the number of factors with the hidden size $\mathcal{O}(n_{e}n_{v}+L)$, if the Transformer can learn a perfect weights component in layer $k$, then it can learn a representation given $i$ in-context examples with context comp. score $\operatorname{srs}_{i}=(1-\operatorname{srs}_{i-1})s_{i}+\operatorname{srs}_{i-1}$ and $\operatorname{srs}_{0}=s_{0}$ at layer $k+2$, where $s_{i}=1-\sum_{j=0}^{i}\tbinom{i}{j}\sum_{k=2}^{|E|}\tbinom{|E|}{k}\frac{k-1}{k}\left(\frac{(n_{v}-1)^{i-j}}{n_{v}^{i}}\right)^{k}\left(1-\frac{(n_{v}-1)^{i-j}}{n_{v}^{i}}\right)^{|E|-k}$, and we can obtain the accuracy as $\operatorname{cls}_{i}=\frac{(n_{v}-1)(n_{v}^{i-1}-(n_{v}-1)^{i-1})}{n_{v}^{i}}\operatorname{srs}_{i}+\frac{1}{n_{v}}$ at $k+3$ layers.
The constucted Transformer achieve significant performance.  To illustrate that the construction is meaningful, we compare the performance of the constructed model with the previously trained model. We choose the performance of the model trained on $D_{\text{fix}\land\text{rnd}}$ settings as it is the best-performed model on $D_{\text{rnd}}$ setting. We find that the constructed model achieved a comparable performance as that of the trained model. These results indicate that our construction is meaningful.
@@ E.1 Proof Sketch
1. Estimate the hidden factor.  According to the perfect weights component assumption, for any $j$, we can project the $j$-th token feature into the space ${\mathbf{f}}^{(e)}_{j}$. Assuming the $j$-th token is not the prompt token, then we have its label information $y_{j}$. Then, for $i$-th token, where $i<j$, obviously $i$ is also not the prompt token. As a result, we also have the information about $y_{i}$. If a factor $e$ is the hidden factor, then we would expect ${\mathbf{y}}_{i}\cdot{\mathbf{y}}_{j}={\mathbf{m}}(v^{(e_{h})}_{i})\cdot{\mathbf{m}}(v^{(e_{h})}_{j})={\mathbf{v}}^{(e_{h})}_{i}\cdot{\mathbf{v}}^{(e_{h})}_{j}={\mathbf{f}}_{i}^{(e_{h})}\cdot{\mathbf{f}}_{j}^{(e_{h})}$, where ${\mathbf{v}},{\mathbf{y}},{\mathbf{m}}$ is the corresponding one-hot version of $v,y,m$. Therefore, if we can find $e$ such that ${\mathbf{f}}_{i}^{(e)}\cdot{\mathbf{f}}_{j}^{(e)}$ can have a same value as ${\mathbf{y}}_{i}\cdot{\mathbf{y}}_{j}$, for all $i$, then $e$ can be predicted as hidden factor. Based on this intuition, in the construction, we focus on finding a way to compare the value between ${\mathbf{f}}_{i}^{(e)}\cdot{\mathbf{f}}_{j}^{(e)}$ and ${\mathbf{y}}_{i}\cdot{\mathbf{y}}_{j}$. All these operations are done in the first layer.
2. Block unrelated information.  After finding the hidden factor, the next step is to block the information that is unrelated to the hidden factor. Blocking unrelated information can remove the influence of it and simplify the following steps. We add a large negative value to the positions of the representation that is unrelated to the hidden factor. Then, through the Relu operation, all these negative values will be removed. These operations are placed in Layer 2.
3. Predict $y$.  The final step is to predict the $y_{p}$ for the prompt sample ${\mathbf{x}}_{p}$. The challenge here is that we don’t know the mapping function $m$ that bridges the factor value and the corresponding label. Consider the relation that for any $i,j$ satisfying $i<j$, we have ${\mathbf{v}}_{i}^{(e_{h})}\cdot{\mathbf{v}}_{j}^{(e_{h})}={\mathbf{f}}_{i}^{(e_{h})}\cdot{\mathbf{f}}_{j}^{(e_{h})}={\mathbf{y}}_{i}\cdot{\mathbf{y}}_{j}$. Therefore, if we can find a $i$, satisfying that ${\mathbf{f}}_{i}^{(e_{h})}\cdot{\mathbf{f}}_{j}^{(e_{h})}=1$, then we have $y_{i}=y_{j}$. Then, we can copy $y_{i}$ to the representation of $j$-th token and use the final linear layer to output $y_{i}$ as the prediction for $y_{j}$.
@@
Given the perfect weights component assumption, there are two source that cause the error prediction. We give the separate analysis below.
1. Fail to find the correct hidden factor. Finding the correct hidden factor is essential to make the correct prediction. The correct hidden factor cannot be inferred if there are more than two factors such that the values of ${\mathbf{f}}_{i}^{(e)}\cdot{\mathbf{f}}_{j}^{(e)}$ and ${\mathbf{y}}_{i}\cdot{\mathbf{y}}_{j}$ can be matched according to the construction of first layer.
2. Make error prediction and give the correct factor. Given the hidden factor, it is also possible that we cannot make the correct prediction if we fail to associate the correct label the factor value. This will happen if all the context examples don’t contain the same factor value of the hidden factor as the prompt image.
@@ E.3 Construction of Transformer
Without loss of generality, we assume the representation of the Transformer in layer $k$ is in a form that ${\mathbf{h}}_{2i-1}^{(k)}=({\mathbf{f}}_{i},\mathbf{0},\mathbf{0},\mathbf{0},\mathbf{0},\mathbf{0},\mathbf{0},{\mathbf{p}}_{i})^{\mathrm{T}}$ and ${\mathbf{h}}_{2i}^{(k)}=(\mathbf{0},{\mathbf{y}}_{i},\mathbf{0},\mathbf{0},\mathbf{0},\mathbf{0},\mathbf{0},{\mathbf{p}}_{i})^{\mathrm{T}}$ (Remind that one sample will take two token, one for $x$ and one for $y$). Because the representation usually lies in low-dimension space, a simple linear layer can transfer the representation in our defined sparse form. Moreover, it is natural to assume that the position information is stored in the representation since it is given in the input and is essential for attention.
@@
This layer obtains the logit of the new sample by comparing the similarity between the unblocked feature of this sample and the in-context sample.
@@
Note that value ${\mathbf{f}}_{i,x}^{\prime}\cdot{\mathbf{f}}_{a,y}^{\prime}$ is equal to the number of unblocked factors (both unblocked) that have the same value between $a$-th sample and $i$-th sample Obviously, there is a $W_{O}^{(l+3,1)}$ such that ${\mathbf{h}}_{2i-1}=({\mathbf{f}}_{i,x}^{\prime},\mathbf{0},{\mathbf{u}}_{2i-1},\mathbf{0},\mathbf{0},\mathbf{0},{\mathbf{m}}_{2i-1}^{\prime},\text{Logit},{\mathbf{p}}_{i})$.